### v2

@@ 4.2 Overview of SoftHGNN
Based on the concept of soft hypergraphs, we propose a novel hypergraph computation framework, namely the Soft Hypergraph Neural Network (SoftHGNN), as shown in Figure 1. This architecture is designed to dynamically generate soft hyperedges that effectively capture high-order semantic relationships in visual data, addressing the limitations of traditional hypergraphs in terms of hyperedge redundancy and rigid binary associations. Furthermore, SoftHGNN is a plug-and-play module that can be easily inserted into any model for visual recognition tasks and significantly improves the performance of the model. Overall, SoftHGNN consists of the following three core parts:
@@ 4.3 Soft Hyperedge Generation
The soft hyperedge generation module dynamically constructs a continuous and differentiable participation matrix $A$ based on the input vertex features, enabling each vertex to participate in each hyperedge to varying degrees. This module not only leverages local vertex features but also fully utilizes global contextual information. By incorporating multi-head mechanisms and pre-projection strategies, it achieves fine-grained capture of high-order semantic relationships. The generation of soft hyperedges comprises three components: dynamic hyperedge prototype generation, vertex feature pre-projection, and participation matrix generation.
@@ 4.3.1 Dynamic Hyperedge Prototype Generation
Specifically, let the vertex feature matrix of a sample be $X\in\mathbb{R}^{N\times D}$, where $N$ is the number of vertices (i.e., tokens). For this sample, we denote the feature representations of the vertices as $\{x_{1},x_{2},\ldots,x_{N}\},\quad x_{i}\in\mathbb{R}^{D}.$ To extract global contextual information, we perform average pooling over all vertices to obtain $f_{\text{avg}}$, and we perform max pooling over each feature dimension to obtain the max-pooled feature $f_{\text{max}}$ By concatenating these two pooled results along the feature dimension, we obtain the global context vector for the sample: $f_{\text{global}}=\begin{pmatrix}f_{\text{avg}},f_{\text{max}}\end{pmatrix}\in\mathbb{R}^{2D}$. Next, a context-aware offset prediction network $\phi$ processes the global context vector to generate the hyperedge prototype offset:
@@ 4.3.2 Vertex Feature Pre-projection
To enhance semantic representation, we pre-project the vertex features and introduce a multi-head mechanism. Specifically, the input vertex features are mapped to a new space via a linear transformation:
@@
where $W_{\text{pre}}\in\mathbb{R}^{D\times D}$ is the pre-projection parameter matrix. Subsequently, the pre-projected vertex features are split into $h$ heads, with each head having a dimension of $D_{\text{head}}=\frac{D}{h}$. We denote the multi-head vertex features as $X_{\text{heads}}\in\mathbb{R}^{h\times N\times D_{\text{head}}}.$ Similarly, the dynamic hyperedge prototypes $P\in\mathbb{R}^{M\times D}$ are divided into heads, yielding $P_{\text{heads}}\in\mathbb{R}^{h\times M\times D_{\text{head}}}.$ The multi-head partition allows the model to capture a more diverse set of semantic relationships in different subspaces, thereby enhancing the expressive capacity of the hyperedge generation.
@@ 4.3.3 Participation Matrix Generation
After obtaining the multi-head representations of vertex features and hyperedge prototypes, we compute the participation matrix $A$ by measuring the dot-product similarity, which quantifies the degree of participation of each vertex in each hyperedge. For each head $\tau\in\{1,2,\dots,h\}$, the similarity between vertices and hyperedges is computed as:
@@
Here, dividing by $\sqrt{D_{\text{head}}}$ stabilizes the gradients by preventing the dot-product values from growing too large as the dimensionality increases. To fuse the information across all heads, we average the similarity scores: $S=\frac{1}{h}\sum_{\tau=1}^{h}S^{(\tau)}\in\mathbb{R}^{N\times M}$. Finally, we apply Softmax along the node dimension for each hyperedge’s node assignment vector, yielding the continuous and differentiable participation matrix:
@@
where $i\in\{1,2,\ldots,N\}$ and $j\in\{1,2,\ldots,M\}$.
This dot-product similarity computation allows each vertex to participate in each hyperedge to varying degrees. Such a soft allocation mechanism enables the model to finely adjust the contribution of each vertex to the hyperedges, especially in complex semantic scenarios where ambiguity or partial overlap may occur.
@@ 4.5 Sparse Soft Hyperedge Selection and Load-Balancing
Typically, the basic SoftHGNN architecture is sufficient for most scenarios. However, in more complex situations (e.g., dense crowds, high-resolution images), the input data may contain a richer and more diverse set of high-order relationships, necessitating an increase in the number of soft hyperedges. To expand the number of soft hyperedges while avoiding an explosion in computational complexity, and inspired by Mixture-of-Experts (MoE) [37, 38], we propose a Sparse Hyperedge Selection (SeS) strategy, combined with load-balancing regularization. This strategy is based on the following hypothesis: In complex scenes, although there exist a greater number of and more diverse high-order relationships, only a small portion of these relationships are crucial for making accurate recognition or prediction decisions. Therefore, we need to employ a larger set of hyperedges to capture the richer high-order associations, but only select the most important ones for message passing and feature updating to avoid redundant computation. Furthermore, to prevent a few hyperedges from being over-activated or under-utilized, we introduce a load-balancing regularizer to encourage all hyperedges to contribute uniformly during training. In the following, we describe the sparse hyperedge selection strategy and the load-balancing regularizer, respectively.
@@ 4.5.1 Sparse Hyperedge Selection Strategy
In the soft hyperedge generation module, assume that $M$ hyperedges are obtained, with the initial (unnormalized) participation matrix given by $S\in\mathbb{R}^{N\times M}$. We partition the soft hyperedge set $\mathcal{E}$ into two subsets:
@@
where $S^{\text{dyn}}$ denotes the portion of the initial participation matrix $S$ corresponding to the $M_{\text{dyn}}$ dynamic soft hyperedges. For the dynamic soft hyperedges, we select only the top-$k$ hyperedges with the highest activation scores. Theoretically, this step can be viewed as selecting a subset $\mathcal{E}^{\text{sel}}\subset\mathcal{E}^{\text{dyn}}$ with $|\mathcal{E}^{\text{sel}}|=k$. Finally, the hyperedge set is updated as $\mathcal{E}=\mathcal{E}^{\text{fixed}}\cup\mathcal{E}^{\text{sel}},$ and the corresponding updated unnormalized participation matrix becomes $S^{\prime}\in\mathbb{R}^{N\times(M_{\text{fixed}}+k)}.$ After applying Softmax to each hyperedge’s node assignment vector, we obtain the final continuous and differentiable participation matrix used for message passing:
@@ 4.5.2 Load-Balancing Regularizer
To avoid situations where some soft hyperedges are over-activated or remain largely inactive, we introduce a load-balancing regularizer. During training, we record the selection outcomes of the dynamic soft hyperedges over multiple forward passes. Specifically, suppose that over $T$ forward passes we construct a binary mask $\mathcal{M}\in\mathbb{R}^{T\times M_{\text{dyn}}}$, where each entry is defined as
@@
This regularization term encourages the dynamic soft hyperedges to be selected uniformly over multiple forward passes, ensuring that each dynamic hyperedge contributes meaningfully to the message passing process rather than relying solely on a few hyperedges, thereby enhancing the model’s robustness as well as the effective utilization of hyperedges.
@@ 4.6.1 The Stage of Soft Hyperedge Generation
Multi-Head Projection and Similarity Calculation. After mapping the vertex features and hyperedge prototypes into a multi-head subspace, the dot-product similarity between vertices and the hyperedge prototype is computed. The dot-product similarity calculation for vertices and the prototype requires about $\mathcal{O}(N\times M\times(D/h))$ multiplications per head, and after combining all heads, the complexity is $\mathcal{O}(h\times N\times M\times(D/h))=\mathcal{O}(N\times M\times D)$.
Thus, the main computational cost in the soft hyperedge generation stage arises from the multi-head projection and similarity calculation, and overall the time complexity for this stage is approximately $\mathcal{O}(N\times M\times D)$.
@@ 5.1 Tasks and Datasets
To validate the superiority and universality of the proposed SoftHGNN, we conducted extensive experiments on three typical visual recognition tasks: image classification, crowd counting, and object detection. In these tasks, we employed a total of five widely-used datasets. In the following, we briefly introduce these datasets by task.
Image Classification. For the image classification task, we adopted the CIFAR-10 [16] and CIFAR-100 [16] datasets, which are widely used in both academia and industry to evaluate model performance on color image classification. The CIFAR-10 dataset consists of 60,000 images divided into 10 classes. Each class contains 6,000 images, with 5,000 images for training and 1,000 images for testing. The CIFAR-100 is an extension of CIFAR-10, containing 60,000 images divided into 100 classes. Each class comprises 600 images, with 500 images for training and 100 images for testing.
Crowd Counting. The crowd counting task focuses on accurately estimating the number of targets (persons) in dense scenes. We utilized the ShanghaiTech Part-A and Part-B datasets [17]. In each image, every target is annotated with a point near the center of the head, resulting in over 330,000 annotations in total. ShanghaiTech Part-A: contains 482 images, primarily featuring dense crowd scenes, with 300 images in the training set and 182 images in the test set. ShanghaiTech Part-B comprises 716 images with relatively sparser crowd scenes, where 400 images are used for training and 316 images for testing.
Object Detection. The object detection task requires the model to not only recognize the categories of objects in images but also to accurately localize them. For this purpose, we conducted experiments on the MS COCO dataset [18]. The MS COCO dataset is a large-scale benchmark for object detection, containing over 328,000 images. It provides bounding boxes and instance segmentation masks for 80 object categories, along with extensive keypoint annotations, making it an essential benchmark in the field of object detection.
@@ 5.2 Evaluation Metrics
For the image classification task, we use accuracy as the primary evaluation metric.
In the crowd counting task, we use Mean Absolute Error (MAE) and Mean Squared Error (MSE) as evaluation metrics, following previous studies [39, 40, 41].
@@
Table 1: Classification accuracy (%) on CIFAR-10 and CIFAR-100 datasets.
Method CIFAR-10 CIFAR-100
ViT [4] 78.6 53.4
+ Self-Attention [4] 78.8 ($\uparrow$0.2) 53.8 ($\uparrow$0.4)
+ HGNN [22] 78.5 ($\downarrow$0.1) 53.6 ($\uparrow$0.2)
+ SoftHGNN 80.5 ($\uparrow$1.9) 55.2 ($\uparrow$1.8)
+ SoftHGNN-SeS 81.1 ($\uparrow$2.5) 55.8 ($\uparrow$2.4)
@@ 5.3 Experimental Setup
CSRNet [42] VGG-16 [43] 68.2 115.0 10.6 16.0 16.3 M
BL [44] VGG-19 [43] 62.8 101.8 7.7 12.7 21.5 M
DM-Count [45] VGG-19 [43] 59.7 95.7 7.4 11.8 21.5 M
TopoCount [46] VGG-16 [46] 61.2 104.6 7.8 13.7 25.8 M
LSC-CNN [41] VGG-16 [43] 66.4 117.0 8.1 12.7 35.1 M
Ctrans-MISN [47] Swin-Trans [6] 55.8 95.9 7.3 11.4 -
CLTR [48] ViT [4] 56.9 95.2 6.5 10.6 43.5 M
GMS [49] HRNet [50] 68.8 138.6 16.0 33.5 -
GCFL [51] VGG-19 [43] 57.5 94.3 6.9 11.0 21.5 M
Gramformer [52] VGG-19 [43] 54.7 87.1 - - 29.1 M
CAAPN [40] VGG-16 [43] 54.4 97.3 5.8 9.8 -
mPrompt [53] HRNet [50] 53.2 85.4 6.3 9.8 48.3 M
CCTrans [39] Twins-PCPVT [54] 54.8 89.0 7.6 12.2 106.4 M
+ Self-Attn [4] Twins-PCPVT [54] 54.7 ($\downarrow$0.2%) 89.7 ($\uparrow$0.8%) 7.4 ($\downarrow$2.6%) 10.9 ($\downarrow$10.7%) 107.6 M
+ HGNN [22] Twins-PCPVT [54] 53.7 ($\downarrow$2.0%) 82.7 ($\downarrow$7.1%) 7.2 ($\downarrow$5.3%) 10.9 ($\downarrow$10.7%) 108.0 M
+ SoftHGNN Twins-PCPVT [54] 52.8 ($\downarrow$3.6%) 82.1 ($\downarrow$7.8%) 7.1 ($\downarrow$6.6%) 10.6 ($\downarrow$13.1%) 107.6 M
+ SoftHGNN-SeS Twins-PCPVT [54] 53.4 ($\downarrow$2.6%) 82.0 ($\downarrow$7.9%) 7.0 ($\downarrow$7.9%) 10.6 ($\downarrow$13.1%) 112.9 M
CCTrans [39] Pyramid ViT [55] 53.7 85.9 7.4 12.5 49.9 M
+ Self-Attn [4] Pyramid ViT [55] 53.1 ($\downarrow$1.1%) 83.1 ($\downarrow$3.3%) 7.1 ($\downarrow$4.1%) 10.9 ($\downarrow$12.8%) 51.2 M
+ HGNN [22] Pyramid ViT [55] 53.5 ($\downarrow$0.4%) 85.5 ($\downarrow$0.5%) 7.0 ($\downarrow$5.4%) 10.9 ($\downarrow$12.8%) 51.6 M
+ SoftHGNN Pyramid ViT [55] 51.7 ($\downarrow$3.7%) 81.3 ($\downarrow$5.4%) 6.8 ($\downarrow$8.1%) 10.7 ($\downarrow$14.4%) 51.2 M
+ SoftHGNN-SeS Pyramid ViT [55] 51.7 ($\downarrow$3.7%) 79.2 ($\downarrow$7.8%) 6.6 ($\downarrow$10.8%) 10.5 ($\downarrow$16.0%) 56.4 M

### v3

@@ 3.3.2 Complexity of Hypergraph Neural Networks
Figure 1: Overview of the Soft Hypergraph Neural Network (SoftHGNN) pipeline. (1) Soft Hyperedge Generation. A set of learnable prototype vectors is dynamically offset by a context‑aware network to form sample‑specific soft hyperedges. Then each vertex computes participation weights to build the continuous participation matrix $A$. (2) Message Passing on Soft Hypergraph. High‑order interactions are captured via a two‑stage aggregation from vertices to hyperedges ($\mathcal{V}\!\to\!\mathcal{E}$) and from hyperedges to vertices ($\mathcal{E}\!\to\!\mathcal{V}$). (3) Sparse Hyperedge Selection (optional). The sparse hyperedge selection strategy is introduced to expand the number of hyperedges while avoiding a significant increase in computational complexity.
@@ 4 Proposed Method
In this section, we present the proposed SoftHGNN in six parts. Section 4.1 gives the definition of a Soft hypergraph. Section 4.2 provides an overall introduction to the SoftHGNN workflow. Section 4.3 and Section 4.4 describe the construction of the soft hypergraph and the message passing on the soft hypergraph, respectively. Section 4.5 introduces sparse soft hyperedges and the load-balancing regularization algorithm. Section 4.6 discusses the computational complexity of SoftHGNN. Table 1 lists the main symbols and their descriptions for easy reference.
@@ 4.2 Overview of SoftHGNN
Based on the concept of soft hypergraphs, we propose a novel hypergraph computation framework, namely the Soft Hypergraph Neural Network (SoftHGNN), as shown in Figure 1. This architecture is designed to dynamically generate soft hyperedges that effectively capture high-order semantic relationships in visual data, addressing the limitations of traditional hypergraphs in terms of hyperedge redundancy and rigid binary associations. Furthermore, SoftHGNN is a lightweight add-on module that can be attached to semantically mature stages of existing vision pipelines for visual recognition tasks. Overall, SoftHGNN consists of the following three core parts:
@@ 4.3 Soft Hyperedge Generation
The soft hyperedge generation module dynamically constructs a continuous and differentiable participation matrix $A$ based on the input vertex features, enabling each vertex to participate in each hyperedge to varying degrees. This module not only leverages local vertex features but also fully utilizes global contextual information. By incorporating a head-wise partition mechanism and pre-projection strategy, it achieves fine-grained capture of high-order semantic relationships. The generation of soft hyperedges comprises three components: dynamic hyperedge prototype generation, vertex feature pre-projection, and participation matrix generation.
@@ 4.3.1 Dynamic Hyperedge Prototype Generation
Specifically, let the vertex feature matrix of a sample be $X\in\mathbb{R}^{N\times D}$, where $N$ is the number of vertices (i.e., tokens). For this sample, we denote the feature representations of the vertices as $\{x_{1},x_{2},\ldots,x_{N}\},\quad x_{i}\in\mathbb{R}^{D}.$ To extract global contextual information, we perform average pooling over all vertices to obtain $f_{\text{avg}}$, and we perform max pooling over each feature dimension to obtain the max-pooled feature $f_{\text{max}}$. By concatenating these two pooled results along the feature dimension, we obtain the global context vector for the sample: $f_{\text{global}}=\begin{pmatrix}f_{\text{avg}},f_{\text{max}}\end{pmatrix}\in\mathbb{R}^{2D}$. Next, a context-aware offset prediction network $\phi$ processes the global context vector to generate the hyperedge prototype offset:
@@ 4.3.2 Vertex Feature Pre-projection
To enhance semantic representation, we pre-project the vertex features and introduce a head-wise partition mechanism. Specifically, the input vertex features are mapped to a new space via a linear transformation:
@@
where $W_{\text{pre}}\in\mathbb{R}^{D\times D}$ is the pre-projection parameter matrix. Subsequently, the pre-projected vertex features are split into $h$ heads, with each head having a dimension of $D_{\text{head}}=\frac{D}{h}$. We denote the head-wise vertex features as $X_{\text{heads}}\in\mathbb{R}^{h\times N\times D_{\text{head}}}.$ Similarly, the dynamic hyperedge prototypes $P\in\mathbb{R}^{M\times D}$ are divided into heads, yielding $P_{\text{heads}}\in\mathbb{R}^{h\times M\times D_{\text{head}}}.$ This head-wise partition decomposes the similarity computation into multiple lower-dimensional groups and then aggregates the resulting scores, reducing reliance on a single full-dimensional similarity estimate for participation matrix generation.
@@ 4.3.3 Participation Matrix Generation
After obtaining the head-wise representations of vertex features and hyperedge prototypes, we compute the participation matrix $A$ by measuring the dot-product similarity, which quantifies the degree of participation of each vertex in each hyperedge. For each head $\tau\in\{1,2,\dots,h\}$, the similarity between vertices and hyperedges is computed as:
@@
Here, dividing by $\sqrt{D_{\text{head}}}$ stabilizes the gradients by preventing the dot-product values from growing too large as the dimensionality increases. To fuse the information across all heads, we average the similarity scores from each head:
$S=\frac{1}{h}\sum_{\tau=1}^{h}S^{(\tau)}\in\mathbb{R}^{N\times M}.$ (8)
This operation produces a single vertex-hyperedge score matrix for the generation of the participation matrix $A$.
Finally, we apply Softmax normalization to the similarity score matrix $S$ to obtain the continuous and differentiable participation matrix $A$. Specifically, we provide two normalization approaches. The first is normalization along the hyperedge dimension (referred to as E-Norm), which is our default approach. Its purpose is to ensure that the sum of the participation scores of all vertices within the same hyperedge equals one. Formally:
@@
where $i\in\{1,2,\ldots,N\}$ and $j\in\{1,2,\ldots,M\}$. The second method is normalization along the vertex dimension (referred to as V-Norm). It ensures that for a given vertex, the sum of its participation scores across all hyperedges equals 1. Formally:
$A_{i,j}=\frac{\exp(S_{i,j})}{\sum_{k=1}^{M}\exp(S_{i,k})}.$ (10)
Such a soft allocation mechanism enables the model to finely adjust the contribution of each vertex to the hyperedges, especially in complex semantic scenarios where ambiguity or partial overlap may occur.
@@ 4.5 Sparse Soft Hyperedge Selection and Load-Balancing
Typically, the basic SoftHGNN architecture is sufficient for most scenarios. However, in more complex situations (e.g., dense crowds or high-resolution images), the input may contain a richer and more diverse set of high-order relationships, which calls for a larger candidate hyperedge space. A naive increase in the number of soft hyperedges, however, does not necessarily improve performance: it may introduce overlapping or weakly useful hyperedges, cause semantic diffusion across hyperedges, or lead the model to over-rely on a small subset of hyperedges while leaving others under-utilized.
To address this issue, and inspired by sparsification in Mixture-of-Experts (MoE) [moe_survey, deepseekmoe], we propose a Sparse Hyperedge Selection (SeS) strategy combined with load-balancing regularization. The goal is not only to control computation, but also to encourage specialization among candidate hyperedges when the hyperedge capacity is expanded. Specifically, we employ a larger set of candidate hyperedges to cover richer high-order associations, while sparsely selecting only the most important ones for message passing and feature updating. Meanwhile, the load-balancing regularizer discourages the model from over-activating only a few hyperedges and promotes more balanced utilization across the candidate set. In the following, we describe the sparse hyperedge selection strategy and the load-balancing regularizer, respectively.
@@ 4.5.1 Sparse Hyperedge Selection Strategy
In the soft hyperedge generation module, assume that $M$ hyperedges are obtained, with the initial (unnormalized) participation matrix given by $S\in\mathbb{R}^{N\times M}$. Note that $S$ denotes the unnormalized similarity matrix before applying Softmax (see Section 4.3.3). We partition the soft hyperedge set $\mathcal{E}$ into two subsets:
@@
where $S^{\text{dyn}}$ denotes the portion of the initial participation matrix $S$ corresponding to the $M_{\text{dyn}}$ dynamic soft hyperedges. We then rank the dynamic hyperedges by their global activation scores $\{g_{j}\}$ and select the Top-$k$ ones. Theoretically, this step can be viewed as selecting a subset $\mathcal{E}^{\text{sel}}\subset\mathcal{E}^{\text{dyn}}$ with $|\mathcal{E}^{\text{sel}}|=k$. Finally, the hyperedge set is updated as $\mathcal{E}=\mathcal{E}^{\text{fixed}}\cup\mathcal{E}^{\text{sel}},$ and the updated unnormalized participation matrix is formed by the concatenation of these hyperedges’ scores. As a result, the corresponding updated unnormalized participation matrix becomes $S^{\prime}\in\mathbb{R}^{N\times(M_{\text{fixed}}+k)}.$ Then, by applying Softmax normalization to the $S^{\prime}$, we obtain the final continuous and differentiable participation matrix used for message passing:
@@ 4.5.2 Load-Balancing Regularizer
To avoid situations where some soft hyperedges are over-activated or remain largely inactive, we introduce a load-balancing regularizer.
This design is analogous to the auxiliary load-balancing losses used in MoE-style sparse activation to prevent expert collapse, but here it is applied to the selection of dynamic soft hyperedges and avoids unbalanced use of these soft hyperedges. During training, we record the selection outcomes of the dynamic soft hyperedges over multiple forward passes. Specifically, suppose that over $T$ forward passes we construct a binary mask $\mathcal{M}\in\mathbb{R}^{T\times M_{\text{dyn}}}$, where each entry is defined as
@@
This regularization term encourages the dynamic soft hyperedges to be selected more uniformly over multiple forward passes, ensuring that each dynamic hyperedge contributes meaningfully to the message passing process rather than relying disproportionately on only a few hyperedges. Beyond balanced utilization, this also promotes functional specialization among candidate hyperedges and mitigates hyperedge collapse when the candidate hyperedge space is enlarged.
@@ 4.6.1 The Stage of Soft Hyperedge Generation
Head-wise Projection and Similarity Calculation. After mapping the vertex features and hyperedge prototypes into a head-wise subspace, the dot-product similarity between vertices and the hyperedge prototype is computed. The dot-product similarity calculation for vertices and the prototype requires about $\mathcal{O}(N\times M\times(D/h))$ multiplications per head, and after combining all heads, the complexity is $\mathcal{O}(h\times N\times M\times(D/h))=\mathcal{O}(N\times M\times D)$.
Thus, the main computational cost in the soft hyperedge generation stage arises from the head-wise projection and similarity calculation, and overall the time complexity for this stage is approximately $\mathcal{O}(N\times M\times D)$.
@@ 5.1 Tasks and Datasets
To validate the superiority and universality of the proposed SoftHGNN, we conduct extensive experiments on three typical visual recognition tasks: image classification, crowd counting, and object detection. In these tasks, we employ a total of five widely-used datasets. In the following, we briefly introduce these datasets by task.
Image Classification. For the image classification task, we adopt the ImageNet-100 [imgnet] and ImageNet-1k [imgnet] datasets. ImageNet-1k is a widely used large-scale benchmark derived from the ILSVRC 2012 challenge, covering 1,000 diverse object categories. It consists of approximately 1.28 million images for training and 50,000 images for validation. ImageNet-100 is a subset constructed by selecting 100 classes from the original ImageNet-1k dataset. It comprises roughly 130,000 training images and 5,000 validation images, serving as a lighter benchmark to evaluate model performance on a smaller scale.
Crowd Counting. The crowd counting task focuses on accurately estimating the number of targets (persons) in dense scenes. We utilized the ShanghaiTech Part-A and Part-B datasets [shanghaitech]. In each image, every target is annotated with a point near the center of the head, resulting in over 330,000 annotations in total. ShanghaiTech Part-A: contains 482 images, primarily featuring dense crowd scenes, with 300 images in the training set and 182 images in the test set. ShanghaiTech Part-B comprises 716 images with relatively sparser crowd scenes, where 400 images are used for training and 316 images for testing.
Object Detection. The object detection task requires the model to not only recognize the categories of objects in images but also to accurately localize them. For this purpose, we conducted experiments on the MS COCO dataset [mscoco]. The MS COCO dataset is a large-scale benchmark for object detection, containing over 328,000 images. It provides bounding boxes and instance segmentation masks for 80 object categories, along with extensive keypoint annotations, making it an essential benchmark in the field of object detection.
@@ 5.2 Evaluation Metrics
For the image classification task, we use Top-1 and Top-5 accuracy as the primary evaluation metric, following mainstream works [swin, pvt, vig].
In the crowd counting task, we use Mean Absolute Error (MAE) and Mean Squared Error (MSE) as evaluation metrics, following previous studies [cctrans, caapn, lsc_cnn].
@@ 5.3 Experimental Setup
Table 2: Comparison of classification performance (%) with different models on ImageNet-100.
Method Top-1 Top-5 Param.
ResNet-18 [resnet] 80.7 94.7 11.2 M
ViHGNN-Tiny [vihgnn] 85.4 96.9 7.3 M
PVT-Small [pvt] 84.2 96.6 24.0 M
PVT-Tiny 82.7 95.4 12.8 M
+Attention [vit] 83.0 ($\uparrow$0.3) 95.7 ($\uparrow$0.3) 14.6 M
+HGNN($\epsilon$-ball) [hgnnp] 83.1 ($\uparrow$0.4) 96.3 ($\uparrow$0.9) 14.0 M
+HGNN($k$-NN) [hgnnp] 82.9 ($\uparrow$0.2) 95.2 ($\downarrow$0.2) 14.0 M
+HgVT Block [hgvt] 85.1 ($\uparrow$2.4) 96.5 ($\uparrow$1.1) 14.7 M
+SoftHGNN 86.1 ($\uparrow$3.4) 97.2 ($\uparrow$1.8) 14.2 M
+SoftHGNN-SeS 86.8 ($\uparrow$4.1) 98.0 ($\uparrow$2.6) 16.4 M
SwinT-Tiny [swin] 88.0 97.6 27.6 M
+Attention [vit] 89.3 ($\uparrow$1.3) 98.0 ($\uparrow$0.4) 30.0 M
+HGNN($\epsilon$-ball) [hgnnp] 89.2 ($\uparrow$1.2) 98.0 ($\uparrow$0.4) 28.5 M
+HGNN($k$-NN) [hgnnp] 89.0 ($\uparrow$1.0) 98.0 ($\uparrow$0.4) 28.5 M
+HgVT Block [hgvt] 89.4 ($\uparrow$1.4) 97.8 ($\uparrow$0.2) 28.8 M
+SoftHGNN 90.1 ($\uparrow$2.1) 98.3 ($\uparrow$0.7) 28.6 M
+SoftHGNN-SeS 90.9 ($\uparrow$2.9) 98.7 ($\uparrow$1.1) 30.7 M
Experiments are conducted on NVIDIA RTX 4090 GPUs. For the image classification task, we adopt Pyramid Vision Transformer and Swin-Transformer as the baseline methods. On both ImageNet-100 and ImageNet-1k, we set the learning rate to $1\times 10^{-3}$ and use the Adam optimizer [adam] with a cosine annealing learning rate scheduler. Standard data augmentation techniques, including random cropping, random scaling, and horizontal flipping, are applied during training. On ImageNet-100, the batch size is set to 512, and the epoch number is set to 200. On ImageNet-1k, the batch size is set to 1024, and the epoch number is set to 300. For the crowd counting task, we adopt CCTrans [cctrans] as our baseline method, which is a representative Transformer-based crowd counting framework. We strictly follow the original experimental setup of CCTrans to ensure fairness in comparison. For the object detection task, we select state-of-the-art detection models YOLO11 [yolo11] and YOLOv12 [yolov12] as baselines. We faithfully reproduce and evaluate these models under their original experimental settings [yolo11, yolov12] to ensure a fair and rigorous comparison. When testing the latency and FLOPs of the model, we utilized one Tesla T4 GPU with TensorRT FP16, which is the same as the official settings for YOLO11 and YOLOv12.
Unless otherwise specified, all models are initialized from random weights and trained end-to-end. By default, both the number of heads and the number of soft hyperedges in SoftHGNN are set to 8. In the variant with Soft Hyperedge Selection, referred to as SoftHGNN-SeS, the number of fixed soft hyperedges is set to 16, with an additional 32 dynamic soft hyperedges from which a sparse subset is selected. A sparsity ratio of 50% is applied, meaning that 16 out of the 32 dynamic soft hyperedges are activated during each forward pass. When using SoftHGNN-SeS, the final loss is computed as the sum of load-balancing regularization loss and the original loss of the baseline.
Table 3: Comparison of classification performance (%) with different models on ImageNet-1k.
Method Top-1 Top-5 Param.
ResNet-18 [resnet] 70.6 89.7 11.7 M
ResNet-50 [resnet] 75.3 92.2 25.6 M
ViG [vig] 73.9 92.0 7.1 M
ViHGNN [vihgnn] 74.3 92.5 8.2 M
PVT-Tiny [pvt] 74.8 92.3 13.2 M
+Attention [vit] 75.1 ($\uparrow$0.3) 92.6 ($\uparrow$0.3) 15.1 M
+HGNN($\epsilon$-ball) [hgnnp] 75.0 ($\uparrow$0.2) 92.4 ($\uparrow$0.1) 14.5 M
+HGNN($k$-NN) [hgnnp] 74.9 ($\uparrow$0.1) 92.4 ($\uparrow$0.1) 14.5 M
+HgVT Block [hgvt] 75.1 ($\uparrow$0.3) 92.5 ($\uparrow$0.2) 15.2 M
+SoftHGNN 75.6 ($\uparrow$0.8) 92.9 ($\uparrow$0.6) 14.7 M
+SoftHGNN-SeS 75.9 ($\uparrow$1.1) 93.1 ($\uparrow$0.8) 16.9 M
@@
CSRNet [csrnet] VGG-16 [vgg] 68.2 115.0 10.6 16.0 16.3 M
BL [bl] VGG-19 [vgg] 62.8 101.8 7.7 12.7 21.5 M
DM-Count [dmcount] VGG-19 [vgg] 59.7 95.7 7.4 11.8 21.5 M
TopoCount [topocount] VGG-16 [vgg] 61.2 104.6 7.8 13.7 25.8 M
LSC-CNN [lsc_cnn] VGG-16 [vgg] 66.4 117.0 8.1 12.7 35.1 M
Ctrans-MISN [ctrans_misn] Swin-Trans [swin] 55.8 95.9 7.3 11.4 -
CLTR [cltr] ViT [vit] 56.9 95.2 6.5 10.6 43.5 M
GMS [gms] HRNet [hrnet] 68.8 138.6 16.0 33.5 -
GCFL [gcfl] VGG-19 [vgg] 57.5 94.3 6.9 11.0 21.5 M
Gramformer [gramformer] VGG-19 [vgg] 54.7 87.1 - - 29.1 M
CAAPN [caapn] VGG-16 [vgg] 54.4 97.3 5.8 9.8 -
mPrompt [mprompt] HRNet [hrnet] 53.2 85.4 6.3 9.8 48.3 M
CCTrans [cctrans] Twins-PCPVT [twins] 54.8 89.0 7.6 12.2 106.4 M
+ Self-Attn [vit] Twins-PCPVT [twins] 54.7 ($\downarrow$0.2%) 89.7 ($\uparrow$0.8%) 7.4 ($\downarrow$2.6%) 10.9 ($\downarrow$10.7%) 107.6 M
+ HGNN [hgnnp] Twins-PCPVT [twins] 53.7 ($\downarrow$2.0%) 82.7 ($\downarrow$7.1%) 7.2 ($\downarrow$5.3%) 10.9 ($\downarrow$10.7%) 108.0 M
+ SoftHGNN Twins-PCPVT [twins] 52.8 ($\downarrow$3.6%) 82.1 ($\downarrow$7.8%) 7.1 ($\downarrow$6.6%) 10.6 ($\downarrow$13.1%) 107.6 M
+ SoftHGNN-SeS Twins-PCPVT [twins] 53.4 ($\downarrow$2.6%) 82.0 ($\downarrow$7.9%) 7.0 ($\downarrow$7.9%) 10.6 ($\downarrow$13.1%) 112.9 M
CCTrans [cctrans] Pyramid ViT [pvtv2] 53.7 85.9 7.4 12.5 49.9 M
+ Self-Attn [vit] Pyramid ViT [pvtv2] 53.1 ($\downarrow$1.1%) 83.1 ($\downarrow$3.3%) 7.1 ($\downarrow$4.1%) 10.9 ($\downarrow$12.8%) 51.2 M
+ HGNN [hgnnp] Pyramid ViT [pvtv2] 53.5 ($\downarrow$0.4%) 85.5 ($\downarrow$0.5%) 7.0 ($\downarrow$5.4%) 10.9 ($\downarrow$12.8%) 51.6 M
+ SoftHGNN Pyramid ViT [pvtv2] 51.7 ($\downarrow$3.7%) 81.3 ($\downarrow$5.4%) 6.8 ($\downarrow$8.1%) 10.7 ($\downarrow$14.4%) 51.2 M
+ SoftHGNN-SeS Pyramid ViT [pvtv2] 51.7 ($\downarrow$3.7%) 79.2 ($\downarrow$7.8%) 6.6 ($\downarrow$10.8%) 10.5 ($\downarrow$16.0%) 56.4 M